Calculate the market equilibrium level of price and quantity
Qd = 5000-6P
Qs= 1500+P
1
Expert's answer
2019-05-13T09:27:00-0400

Solution: Demand equation, "Q_d = 5000 - 6P" .......................(1)


Supply equation, "Q_s = 1500 + P" .........................(2)


In equilibrium, "Q_d = Q_s"

or, "5000 - 6P = 1500 + P"

or, "P = 500"


So, Equilibrium quantity , "Q = Q_d = Q_s = 1500 + 500 = 2000" [As, P = 500 unit.]



Answer: Equilibrium price is 500 unit and quantity is 2000 unit.